OpenStudy (anonymous):

consider limx->2 f(x), where f(x)=4x-1 show that |f(x)-7|<4delta if 0<|x-2|

OpenStudy (anonymous):

I'll see what I can do for you. You may have to insert some nuances that I am likely to leave out having not rigorously proved a limit in several years. Ok here goes:\[\left| f(x)-7 \right|=\left| (4x-1)-7 \right|=\left| 4x-8 \right|=4\left| x-2 \right|\] So\[\left| f(x)-7 \right|=4\left| x-2 \right|<4\delta\] from this we can tell that\[0<\left| x-2 \right|<\delta\] as desired.

OpenStudy (anonymous):

I guess actually it should be from\[0<\left| x-2 \right|<\delta\] we can see that the limit exists.. lol. Whatever. I'm pretty sure the math is right anyway. This is what happens when you don't take a mathematical reasoning or proofs class.

OpenStudy (anonymous):

ok let try to figure out your steps

OpenStudy (anonymous):

so I did get that part right in the work I did can you help me with part b

OpenStudy (anonymous):

What's the part b?